OpenStudy (anonymous):

The specific gravity of iron is 7.87, and the density of water at 4.00 ∘C is 1.00 g/cm3. You can use this information to find the density of iron. Find the volume occupied by 8.00g of iron.

OpenStudy (shinalcantara):

\[\ \gamma = \frac{ m }{ V }\] \[G = \frac{ \gamma _{m} }{ \gamma _{w} }\]

OpenStudy (shinalcantara):

since specific gravity is the density of the material over the density of water then you'll have the density of iron as 1.00(7.87) = 7.87g/cm3 the volume is given by the formula: \[\rho = \frac{ m }{ V }\] so volume would be \[V = \frac{ m }{ \rho }\]

OpenStudy (shinalcantara):

\[V = \frac{ 8g }{ 7.87g/cm3 }\]

OpenStudy (shinalcantara):

@sippara recall the formulas for density and specific gravity that's all you need and you're done

OpenStudy (shinalcantara):

this one's hydraulics and geotechnical though o.o

OpenStudy (anonymous):

Thank you